OpenStudy (anonymous):

solve the equation: (y-3)(y-2)=0

OpenStudy (anonymous):

y=2 and y=3

OpenStudy (anonymous):

If (y-3)(y-2) = 0 then either y-3 = 0 or y-2 = 0. Therefore y must be either 3, or 2.

OpenStudy (anonymous):

thank you